# Relativistic Jets and Winds in Radio-Identified Supermassive Black Hole Binary Candidates

### v1

@@ 3.2.1 Model Application
We apply our model to PKS J0805 with a similar set of parameters as PKS 2131: $\theta_{j}=0.1$, $\Gamma_{j}=10$, $\gamma_{e,\min}=10$, $\gamma_{e,\max}=300$, $p=1.2$, $\beta_{w}=0.9$, $\beta_{H}=0.02$, and $r_{\max}=r_{\min}+60\,(\nu/15\text{ GHz})^{-1}$ ly. We use lower jet power $L_{e}+L_{m}=5\times 10^{45}$ erg s-1 and suppose $i=5^{\circ}$, as there is no VLBI estimated viewing inclination. We do not need to adjust the flux by a constant offset to match the OVRO data, but an additional constant $\sim 0.07$ Jy flux is needed to match the ACT data in each band. We show our model results in Fig. 6. The peaks have large variability, so we choose to match the lower luminosity peaks. Our model reproduces the phase shifts in the ACT data well (right panel of Fig. 6). PKS J0805 light curves at more observing frequencies would be needed to validate the trend and place stronger constraints on these parameters.
@@ 4.1 Emission Lines
Figure 7: A schematic of the wind-jet complex from the primary spinning black hole. Line emitting clouds travel in the wind along helical magnetic field lines from the disk in the black hole frame and are possibly squished due to negative velocity gradient.
@@
We propose a model for the broad-line emission variation. In this picture (illustrated in Fig. 7), the broad line region is composed of cold, dense clouds injected into the wind by the disk (Emmering et al., 1992). We envision clouds following helical magnetic field lines in the wind. In the black hole frame, the clouds have velocity
@@
For simplicity, we imagine all the clouds to be at fixed height $r_{0}$ above the black hole, although it is likely that in the real system they will originate from a range of radii. For this illustrative model, we suppose the clouds are distributed as
$f(\rho)=2\frac{\rho}{\rho_{0}^{2}}\exp\left(-\frac{\rho^{2}}{\rho_{0}^{2}}\right)$ (33)
in cylindrical radius and uniformly in $\Phi$. Thus, $v_{cl,0}$ roughly corresponds to the average cloud speed. These choices are merely demonstrative and can easily be modified.
@@
Lines emitted by the distribution of clouds will have central wavelength
$\frac{\lambda(t)}{\lambda_{0}}=\frac{1}{2\pi}\int\left(1-\frac{\vec{v}_{cl}^{\prime}\left(\rho,\Phi,t\right)}{c}\cdot\vec{n}\right)f(\rho)\text{d}\rho\text{d}\Phi,$ (35)
and line width
@@
We show the variation in central wavelength and width for a line with nominal wavelength $\lambda_{0}=4400$ $\mathrm{\AA }$, corresponding to the semi-forbidden $1909$ $\mathrm{\AA }$ CIII] line at $z=1.3$ in Fig. 8. The line centers vary by $\sim 2-5$ $\mathrm{\AA }$ at the nominal $i=3.8^{\circ}$ for PKS 2131 depending on the orbital velocity, while the variation increases for higher inclinations. These small shifts can be measured with several observations per orbit over the course of multiple orbits. Together with the amplitude of the flux variation, these shifts reveal the orbital speed and consequently the component masses. This simple example is meant to be illustrative. The real variability will be more complicated. Only observations will tell.
@@ 4.2 Gravitational Waves
Given their few year periods, PKS 2131 and PKS J0805 likely contribute to the PTA-detected GW background. The GW amplitude and inspiral time depend on the black hole masses, which are at present unknown. $M_{1}\sim 10^{6}-10^{9}$ M give GW characteristic strains $\sim 10^{-17}-10^{-14}$. SMBHBs with $M_{1}\approx 10^{8}$ M represent the lightest systems detectable by the current international network of PTAs at these redshifts, and correspond to our fiducial $\beta_{H}\approx 0.02$. These sources are expected to merge in $10^{4}-10^{5}$ yr, while $10^{7}$ $M_{\odot}$ sources (corresponding to $\beta_{H}\approx 0.01$) will likely merge in $10^{6}-10^{7}$ yr. PKS 2131 and PKS J0805 should both merge well within the age of the universe, but are far enough from merger to remain continuous non-evolving GW sources. These and future radio-identified binaries provide definitive locations on the sky and GW periods, as well as binary phase, orbital velocity, and observing inclination constraints — all important priors for targeted GW searches in PTA data. This motivates searches through the radio PTA data as well as available Fermi gamma-ray pulsar timing data, which can also constrain nHz GW signals (Ajello et al., 2022). NANOGrav PTA data has begun to place upper limits on these systems’ chirp masses ($\lesssim 10^{10}$ M) (Agarwal et al., 2025) based on their sky locations and periods. These searches can be enhanced by adding more priors from our modeling as well as longer observing baselines. Convincing evidence of GWs from even one source would represent a major discovery.
@@ 5 Conclusions
Of the two sources, PKS 2131 remains the stronger SMBHB candidate. Recent ACT observations have strengthened the case for PKS J0805. The few yr periods make these sources some of the most promising individually identifiable nHz GW emitting candidates. Ultimately, more data is required to determine if either is indeed a binary as well as to probe the putative wind. Near-IR observations could bridge the gap between radio and optical to extend the phase shift trend. High quality spectra will be key, as they would contain imprints of the orbital velocities and wind velocities and show variability (e.g. Blundell & Bowler, 2005). A lack of spectral variation would suggest the SMBHB hypothesis is severely deprecated.
Long-term blazar monitoring with OVRO has revealed the existence of this class of possible SMBHBs. Simons Observatory and other mm and cm instruments continuously monitoring the sky may find more sources with similar phenomenology. The periods seen in PKS 2131 and PKS J0805 have been predictive and in PKS 2131 also retrodictive. As PKS 2131 has proven, extant archival data has the ability to make less significant periodicities more significant. Additionally, while the centroid shifts in spatial position associated with these sources may be too small to be resolved by EHT, variation in more nearby sources, which may soon be found, could be directly imaged. Continued monitoring and additional multi-wavelength observing programs can secure these sources as binaries and identify more binary candidates with relativistic jets.
@@ Acknowledgments
The authors thank Alisa Galishnikova, Martijn Oei, Timothy J. Pearson, Viraj Manwadkar, Jack Dinsmore, Roger W. Romani, Emmanuele Sobbachi, Alexander Philippov, Lorenzo Sironi, and Paulo Coppi for useful discussions. This work was supported in part by a grant from the Simons Foundation (00001470, R.B., A.G.S., N.G.). A.G.S. acknowledges the support of the Stanford University Physics Department Fellowship, the National Science Foundation Graduate Research Fellowship, and a Giddings Fellowship at the Kavli Institute for Particle Astrophysics and Cosmology at Stanford. The work at the Owens Valley Radio Observatory is supported by NSF grants AST2407603 and AST2407604. We thank the California Institute of Technology and the Max Planck Institute for Radio Astronomy for supporting the OVRO 40 m program under extremely difficult circumstances over the last 8 years in the absence of agency funding. Without this private support these observations could not have been made. Prior to 2016, the OVRO program was supported by NASA grants NNG06GG1G, NNX08AW31G, NNX11A043G, and NNX13AQ89G from 2006 to 2016 and NSF grants AST-0808050 and AST-1109911 from 2008 to 2014.
Appendix A Towards a more realistic model
The description of orbiting jets, developed in Sec. 2, while sufficient to our current need, ignores jet structure and is limited to small perturbations. We consider additional features for a more detailed analytical model. This may also be useful for interpreting RMHD simulations.
This approach includes three additional features.
  1. 1.
    The velocity of the plasma within the jet is given by Eq. (5), adding the orbital velocity to the radial fluid velocity in the unperturbed jet. The jet has a well-defined surface defined by the jet flow in the barycenter frame.
  2. 2.
    The jet emissivity and absorption coefficient are allowed to vary on spherical surfaces limited by the jet surface. In particular the relativistic electron density can be larger near the jet surface — the sheath — and at its leading edge where it will interact most strongly with the hypothesized confining wind. The radiative transfer is performed on individual rays to compute the observed intensity (Eq. 26) and flux density (Eq. 27).
  3. 3.
    The effective wind-jet boundary velocity $\beta_{w}$ need not be constant with radius. However, its temporal variation is strictly periodic.
Models constructed using this more general approach can help separate the kinematical from the dynamical contributions to the flux variation and will be discussed elsewhere.

### v2

@@ 3.2.1 Model Application
We apply our model to PKS J0805 with a similar set of parameters as PKS 2131: $\theta_{j}=0.1$, $\Gamma_{j}=10$, $\gamma_{e,\min}=10$, $\gamma_{e,\max}=300$, $p=1.2$, $\beta_{w}=0.9$, $\beta_{H}=0.02$, and $r_{\max}=r_{\min}+60\,(\nu(1+z_{S})/15\text{ GHz})^{-1}$ ly. We use lower jet power $L_{e}+L_{m}=5\times 10^{45}$ erg s-1 and suppose $i=5^{\circ}$, as there is no VLBI estimated viewing inclination. We show our model results in Fig. 6. The OVRO peaks have large variability, so we choose to match the lower flux peaks. We do not need to adjust the flux by a constant offset to match the OVRO data, but an additional constant $\sim 0.07$ Jy flux is needed to match the ACT data in each band. The discrepancy in offsets could be due to complexities in emission that we are not modeling. While it is possible to reproduce these features with an ad hoc model, we choose to recover the principal features of the observations with a simple physical description. Note that the ACT light curves correspond to a period where the OVRO flux is high, suggesting that the constant excess needed by the ACT light curves could be the result of a flare in a non-sinusoidally varying emission component. Without much VLBI information for this source, however, we cannot otherwise easily distinguish flux from non-sinusoidal components. Nonetheless, our model reproduces the phase shifts in the ACT data well (right panel of Fig. 6). PKS J0805 light curves at more observing frequencies would be needed to validate the trend and place stronger constraints on these parameters.
@@ 4.1 Emission Lines
Figure 7: A schematic of the wind-jet complex from the primary spinning black hole. Line emitting clouds travel in the wind along helical magnetic field lines from the disk in the black hole frame and are possibly squished due to a negative velocity gradient.
@@
We propose a model for the broad-line emission variation. In this picture (illustrated in Fig. 7), the broad line region is composed of cold, dense clouds injected into the wind by the disk (1992ApJ...385..460E). We envision clouds following helical magnetic field lines in the wind with speeds which are a function of cylindrical radius from the jet axis. In the black hole frame, the clouds have velocity
@@
For simplicity, we imagine the clouds are distributed uniformly in height above the black hole along the $\hat{z}$ axis, non-uniformly in cylindrical radius (in the $x$-$y$ plane), and uniformly in $\Phi$, although it is likely that the real spatial distribution of clouds is more complicated. For this illustrative model, we suppose the clouds are distributed in cylindrical radius as
$f(\rho)=2\frac{\rho}{\rho_{0}^{2}}\exp\left(-\frac{\rho^{2}}{\rho_{0}^{2}}\right).$ (33)
Thus, $v_{cl,0}$ roughly corresponds to the average cloud speed. These choices are merely demonstrative and can easily be modified.
@@
Lines emitted by the distribution of clouds will have central wavelength determined by the average Doppler shift from the distribution of clouds
$\frac{\lambda(t)}{\lambda_{0}}=\frac{1}{2\pi}\int\left(1-\frac{\vec{v}_{cl}^{\prime}\left(\rho,\Phi,t\right)}{c}\cdot\vec{n}\right)f(\rho)\text{d}\rho\text{d}\Phi.$ (35)
The line width will be the standard deviation of the Doppler shifts
@@
We show the variation in central wavelength and width for a line with nominal wavelength $\lambda_{0}=4400$ $\mathrm{\AA }$, corresponding to the semi-forbidden $1909$ $\mathrm{\AA }$ CIII] line at $z_{S}=1.3$ in Fig. 8. The line centers vary by $\sim 2-5$ $\mathrm{\AA }$ at the nominal $i=3.8^{\circ}$ for PKS 2131 depending on the orbital velocity, while the variation increases for higher inclinations. These small shifts can be measured with several observations per orbit over the course of multiple orbits. Together with the amplitude of the flux variation, these shifts reveal the orbital speed and consequently the component masses. This simple example is meant to be illustrative. The real variability will be more complicated. Only more observations will tell.
@@ 4.2 Gravitational Waves
Given their few year periods, PKS 2131 and PKS J0805 likely contribute to the PTA-detected GW background. The GW amplitude and inspiral time depend on the black hole masses, which are at present unknown. $M_{1}\sim 10^{6}-10^{9}$ M give GW characteristic strains $\sim 10^{-17}-10^{-14}$. SMBHBs with $M_{1}\approx 10^{8}$ M represent the lightest systems detectable by the current international network of PTAs at these redshifts, and correspond to our fiducial $\beta_{H}\approx 0.02$. These sources are expected to merge in $10^{4}-10^{5}$ yr, while $10^{7}$ $M_{\odot}$ sources (corresponding to $\beta_{H}\approx 0.01$) will likely merge in $10^{6}-10^{7}$ yr. PKS 2131 and PKS J0805 should both merge well within the age of the universe, but are far enough from merger to remain continuous non-evolving GW sources. These and future radio-identified binaries provide definitive locations on the sky and GW periods, as well as binary phase, orbital velocity, and observing inclination constraints — all important priors for targeted GW searches in PTA data. This motivates searches through the radio PTA data as well as available Fermi gamma-ray pulsar timing data, which can also constrain nHz GW signals (2022Sci...376..521F). NANOGrav PTA data has begun to place upper limits on these systems’ chirp masses ($\lesssim 10^{10}$ M) (2025arXiv250816534A) based on their sky locations and periods. These searches can be enhanced by adding more priors from our modeling as well as longer observing baselines. Convincing evidence of GWs from even one source would represent a major discovery.
@@ 5 Conclusions
This model, while sufficient for our current need, neglects jet structure and is limited to small perturbations. An improved model will need to include three additional features:
  1. 1.
    The jet emissivity and absorption coefficient vary as a function of angular position ($\chi$$\psi$). In particular the relativistic electron density can be larger near the jet surface — the sheath — and at its leading edge where it will interact most strongly with the hypothesized confining wind.
  2. 2.
    The particle acceleration and synchrotron cooling vary as a function of radius along the jet.
  3. 3.
    The effective wind-jet boundary velocity $\beta_{w}$ may vary with radius. However, its temporal variation is strictly periodic.
Models with these features can help separate the kinematical from the dynamical contributions to the flux variation and will be discussed in future work. This may also be useful for interpreting RMHD simulations.
Of the two sources, PKS 2131 remains the stronger SMBHB candidate. Recent ACT observations have strengthened the case for PKS J0805. The few yr periods make these sources some of the most promising individually identifiable nHz GW emitting candidates. Ultimately, more data is required to determine if either is indeed a binary as well as to probe the putative wind. Near-IR observations could bridge the gap between radio and optical to extend the phase shift trend. High quality spectra will be key, as they would contain imprints of the orbital velocities and wind velocities and show variability (e.g. 2005ApJ...622L.129B). A lack of spectral variation would suggest the SMBHB hypothesis is severely deprecated.
Long-term blazar monitoring with OVRO has revealed the existence of this class of possible SMBHBs. Simons Observatory and other mm and cm instruments continuously monitoring the sky may find more sources with similar phenomenology. The periods seen in PKS 2131 and PKS J0805 have been both predictive and retrodictive, as the sinusoidal variation has been corroborated by both new and preexisting data. As PKS 2131 has proven, extant archival data has the ability to make less significant periodicities more significant. Additionally, while the shifts in the spatial position centroid of these sources may be too small to be resolved by EHT, variation in more nearby sources, which may soon be found, could be directly imaged. Continued monitoring and additional multi-wavelength observing programs can secure these sources as binaries and identify more SMBHB candidates with relativistic jets.
@@ Acknowledgments
The authors thank Alisa Galishnikova, Martijn Oei, Timothy J. Pearson, Viraj Manwadkar, Jack Dinsmore, Roger W. Romani, Emmanuele Sobacchi, Alexander Philippov, Lorenzo Sironi, and Paulo Coppi for useful discussions. The authors also thank the anonymous referee for useful feedback which improved this manuscript. This work was supported in part by a grant from the Simons Foundation (00001470, R.B., A.G.S., N.G.). A.G.S. acknowledges the support of the Stanford University Physics Department Fellowship, the National Science Foundation Graduate Research Fellowship, and a Giddings Fellowship at the Kavli Institute for Particle Astrophysics and Cosmology at Stanford. The work at the Owens Valley Radio Observatory is supported by NSF grants AST2407603 and AST2407604. We thank the California Institute of Technology and the Max Planck Institute for Radio Astronomy for supporting the OVRO 40 m program under extremely difficult circumstances over the last 8 years in the absence of agency funding. Without this private support these observations could not have been made. Prior to 2016, the OVRO program was supported by NASA grants NNG06GG1G, NNX08AW31G, NNX11A043G, and NNX13AQ89G from 2006 to 2016 and NSF grants AST-0808050 and AST-1109911 from 2008 to 2014.